OpenStudy (anonymous):

\[give the parametric equestion for the following line.a) the line through (2,9,0) and (-1,5,1) and b)the line through (3,4,5) parallel tp the line x(t)=2-2t , y=4+3t and z=-1-2t

OpenStudy (alexwee123):

lets make A=(2,9,0) and B= (-1,5,1) and O=the origin OA= <2,9,0) and AB=<-3,-4,1> <x,y,z>= <2,9,0> + t<-3,-4,1> so x=2-3t, y=9-4t, and z=t

OpenStudy (alexwee123):

and for x(t)=2-2t , y=4+3t and z=-1-2t you would find the direction vector of this which is <-2,3,-2>

OpenStudy (alexwee123):

and do the same thing i did for the first problem <x,y,z>= <3,4,5> + t<-2,3-2?

OpenStudy (alexwee123):

t<-2,3-2>*

OpenStudy (alexwee123):

so x=3-2t y=4+3t z=5-2t

OpenStudy (alexwee123):

questions? :/

OpenStudy (anonymous):

thank u very much i almost give up but u make my day for c)the line through (1,0,1) perpendicular to the plane 2x-3y+z=9 where does this line intersect the plane

OpenStudy (alexwee123):

well you would first find the perpendicular direction vector of 2x-3y+z=9

OpenStudy (anonymous):

is (2,-3,0)

OpenStudy (anonymous):

it that right

OpenStudy (alexwee123):

well what's the direction vector of 2x-3y+z=9?

OpenStudy (alexwee123):

hint* there are and infinite amount of ways to find the perpendicular vector of a plane because it is 3d

OpenStudy (anonymous):

am still didnt get the right answer i try it many time but i cant answer it right any help please